OpenStudy (anonymous):

Myra’s stamp collection consisted of 120 stamps in October. In March, her collection grew to 138 stamps. Find the percent increase between October and March. wiilll give medal!!!!

OpenStudy (anonymous):

A. 15% B. 9% C. 18% D. 12%

OpenStudy (anonymous):

percent increase formula : percent change = (new value - old value) ------------------- x 100 old value can you take over from here ???

OpenStudy (anonymous):

What @kelliegirl33 posted will work. 120 * 1.x = 138 (I know that's probably worded wrong, but you get it, eh? we need to find what decimal is used to make 120 138, which, since you are given options, you can simply guess and check) Or, Divide the new value, 138, by the old value, 120, to get your answer.

OpenStudy (anonymous):

138 - 120 = 18 18/120 = .15 .15 x 100 = 15% Do you understand how I got to this answer ?